Circular No. 7146

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

η CARINAE

K. Davidson, R. M. Humphreys, and K. Ishibashi, University of Minnesota; T. R. Gull, Goddard Space Flight Center, NASA; M. Hamuy, University of Arizona; and L. Berdnikov and P. Whitelock, South African Astronomical Observatory (SAAO), write: “Hubble Space Telescope STIS observations show that η Car unexpectedly increased in brightness by a factor of ∼ 2 (0.6 to 0.8 mag) between Dec. 1997 and Feb. 1999 at wavelengths around 800 nm. STIS acquisition images, ‘peakup’ countrates (see the STIS instrument handbook), and slit spectra all showed the trend progressively in data obtained on 1997 Dec. 31, 1998 Mar. 19, Nov. 25, and 1999 Feb.

21. These data refer specifically to the central star at spatial resolution ∼ 0′′.1. The images and slit spectra also show similar brightening in at least the inner parts of the surrounding Homunculus nebula, composed of ejecta from the giant eruption seen 150 yr ago. Photometry done at Cerro Tololo and SAAO on 1999 Apr. 17 agree that the Homunculus nebula now has V ∼ 5.25 within a diameter of 20′′, indicating a 0.45-mag difference since the end of 1997 (cf. van Genderen et al. 1999, A.Ap. 343, 847). This is the largest and most rapid brightening of η Car in the past 50 yr, and the object is now brighter than at any time in the past 130 yr. Light-travel-time effects may explain the difference in brightening rates between the star and its surrounding nebula. The phenomenon may continue. Additional photometry and spectroscopy are needed.”

T. S. Metcalfe, University of Texas, Austin; and R. M. Humphreys, University of Minnesota, report: “UBV differential photometry of η Car on Apr. 18 with the 1.5-m reflector (+ Texas 3-channel photometer + 20′′ aperture) at Cerro Tololo Interamerican Observatory gave V = 5.2, B = 5.9, and U = 5.4 (confirming the 0.5-mag rise since Jan. 1998). The aperture included the entire Homunculus nebula.”

NO SUPERNOVA IN UGC 11093 P. Garnavich, S. Jha, P. Challis, and R. Kirshner, Center for Astro-

physics, report that a high-quality spectrum of SN 1999bs was obtained by P. Berlind with the 1.5-m Tillinghast telescope on Apr. 18.5 UT, which indicates that this is not a supernova, as reported on IAUC 7143, but a dwarf nova in outburst. The spectrum now shows rather narrow (1000 km/s halfwidth) Balmer absorption lines from Hβ to Hζ at their rest wavelengths. The low Galactic latitude (+14^{o}) and large separation (nearly 3′) from UGC 11093 is consistent with this object being a Galactic variable.

1999 April 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green